The average salary paid to factory workers is 200 euros. The average salaries paid to secretaries and workers are 260 and 150 euros respectively and the surveillance staff earns 190 euros, they represent 44% of the total workers in the factory. Determine the percentage of secretaries and workers who work in the factory.
Answer #1

here let percentage of secretaries and workers are a and b

therefore a+b+44 =100

or a+b= 56 ............(1)

also average salary =(260*a+150*b+190*44)/100 =200

260a+150b =11640 .....(2)

solving above a = secretaries percentage =29.45 %

and b = workers who work in the factory percentage=26.55 %